# Color3D: Controllable and Consistent 3D Colorization with Personalized Colorizer

### v1

@@ Abstract
In this work, we present Color3D, a highly adaptable framework for colorizing both static and dynamic 3D scenes from monochromatic inputs, delivering visually diverse and chromatically vibrant reconstructions with flexible user-guided control. In contrast to existing methods that focus solely on static scenarios and enforce multi-view consistency by averaging color variations which inevitably sacrifice both chromatic richness and controllability, our approach is able to preserve color diversity and steerability while ensuring cross-view and cross-time consistency. In particular, the core insight of our method is to colorize only a single key view and then fine-tune a personalized colorizer to propagate its color to novel views and time steps. Through personalization, the colorizer learns a scene-specific deterministic color mapping underlying the reference view, enabling it to consistently project corresponding colors to the content in novel views and video frames via its inherent inductive bias. Once trained, the personalized colorizer can be applied to infer consistent chrominance for all other images, enabling direct reconstruction of colorful 3D scenes with a dedicated Lab color space Gaussian splatting representation. The proposed framework ingeniously recasts complicated 3D colorization as a more tractable single image paradigm, allowing seamless integration of arbitrary image colorization models with enhanced flexibility and controllability. Extensive experiments across diverse static and dynamic 3D colorization benchmarks substantiate that our method can deliver more consistent and chromatically rich renderings with precise user control. Project Page https://yecongwan.github.io/Color3D/
@@ 1 Introduction
To mitigate the color inconsistency, a few attempts Dhiman et al. (2023); Cheng et al. (2024b) aims to average multi-view color variations via distillation Dhiman et al. (2023) or dynamic color injection Cheng et al. (2024b). While partially effective, they inevitably dilute the palette richness, producing desaturated and tone-flattened results. Furthermore, smoothing color variations makes the colorized results unpredictable, sacrificing user-controlled colorization ability. Moreover, existing studies only focus on static scenes, while controllable colorization of dynamic ones remains an open and unexplored problem, where color consistency in spatial and temporal dimensions should be maintained.
@@ 3.2 Stage 1: Personalized Colorizer Training
Single View Augmentation. In order to avoid overfitting and more robustly generalize to novel views and video frames, we propose a single view augmentation scheme that combines generative augmentations and traditional augmentations to generate more samples in which the same content consistently retains its color across all instances. As illustrated in Fig. 3(a), the proposed scheme first leverages generative models to produce plausible scene content beyond the provided single view, while ensuring chromatic coherence with the input to enrich the diversity of colorization supervision. Specifically, given a single colored view, we incorporate three complementary generative strategies: (1) Outpainting: we divide the image into a $2\times 2$ grid and apply Stable Diffusion Rombach et al. (2022) to each of the four regions individually to imagine plausible scene content beyond their outer boundaries. Meanwhile, we employ LLaVA Liu et al. (2023a) to generate a descriptive caption based on the image, which serves as the text prompt to guide the diffusion process. (2) Image-to-Video: we generate continuous video frames through Stable Video Diffusion Blattmann et al. (2023) conditioned on the single image to provide coherent object motion and gradual changes in viewpoint. (3) Novel View: to simulate consistent novel viewpoints, we employ Stable Virtual Camera Zhou et al. (2025) to generate views along a predefined orbital trajectory with consistent content across perspectives. Notably, we do not require the generated content to be completely the same as the scene to be colored; instead, we emphasize preserving consistent chromatic styles with the input view, which is sufficient to improve coloring robustness and enhance coloring richness for scene content not included in the single view.
@@ 3.3 Stage 2: 3D Scene Colorization
Table 1: Quantitative comparisons on static (LLFF and Mip-NeRF 360) and dynamic (DyNeRF) 3D scene colorization benchmarks. The top-performing results are highlighted with color.
Method
Language Automatic Reference
@@
GaussianEditor Chen et al. (2024b) 136.49 0.6346 0.093 - - - - - -
GenN2N Liu et al. (2024) - - - 52.63 25.24 0.078 - - -
Ref-NPR Zhang et al. (2023) - - - - - - 111.93 0.7958 0.112
ColorNeRF Cheng et al. (2024b) 138.62 0.6388 0.089 58.26 21.78 0.062 103.68 0.7891 0.086
3DGS+2DColorizer 87.45 0.6415 0.178 40.25 34.07 0.084 67.48 0.7366 0.131
Color3D (Ours) 62.84 0.6544 0.051 35.10 33.99 0.056 41.36 0.6823 0.055
@@
GaussianEditor Chen et al. (2024b) 123.71 0.6045 0.088 - - - - - -
GenN2N Liu et al. (2024) - - - 83.69 23.44 0.137 - - -
Ref-NPR Zhang et al. (2023) - - - - - - 146.32 0.7928 0.143
ColorNeRF Cheng et al. (2024b) 163.12 0.6075 0.093 87.42 18.90 0.093 121.57 0.7852 0.106
3DGS+2DColorizer 112.33 0.6148 0.202 62.32 29.32 0.135 86.25 0.7521 0.178
Color3D (Ours) 68.23 0.6246 0.058 39.03 33.36 0.082 48.62 0.7028 0.079
@@
Instruct 4D-to-4D Mou et al. (2024) 112.35 0.6082 0.062 - - - - - -
4DGS+2DColorizer 89.39 0.6159 0.124 39.17 29.45 0.080 84.65 0.7446 0.128
Color3D (Ours) 58.62 0.6271 0.041 37.28 32.75 0.062 52.77 0.6717 0.052
@@ 4.1 Experimental Settings
Implementation details. All experiments are conducted using the PyTorch framework on NVIDIA RTX A6000 GPUs. We evaluate our method on two static datasets (LLFF Mildenhall et al. (2019) and Mip-NeRF 360 Barron et al. (2022)) and one dynamic dataset (DyNeRF Li et al. (2022)), all of which are converted to monochrome images for colorization. Structure-from-Motion (SfM) is employed to initialize Gaussian points from monochrome inputs in all experiments. Without loss of generality, both our method and competing approaches deploy the same set of image colorization models for fair comparison: ControlColor Liang et al. (2024) for language-guided colorization, DDColor Kang et al. (2023) for automatic colorization, and UniColor Huang et al. (2022) for reference-based colorization. Apart from our proposed Lab Gaussian representation and optimization objectives, all other settings are kept identical to the original 3DGS and 4DGS frameworks. Notably, our entir pipeline incurs only about eight additional minutes for personalized colorizer tuning, which is considered acceptable for chromatic 3D scene reconstruction.
Metrics. We adopt CLIP score Radford et al. (2021) for language-guided colorization to measure text-image alignment, Colorful Score Hasler & Suesstrunk (2003) for automatic colorization to assess color vividness, and Ref-LPIPS Zhang et al. (2023) for reference-based colorization to evaluate perceptual similarity. Fréchet Inception Distance (FID) Heusel et al. (2017) is employed as a standard metric for assessing the rendering quality of all colorization tasks. Additionally, to evaluate the consistency of colorization across views and time, we propose a metric termed Matching Error (ME). It utilizes a dense image matching model Shen et al. (2024) to identify pixel correspondences between ground-truth color images from different views or time steps, and computes the average color difference at these matched locations in the rendered colorized results.
@@ 4.2 Results on Static 3D Scene Colorization
The experimental results for quantitative comparisons in static 3D scenes are shown in the middle and upper portions of Tab. 1. It is observed that our Color3D delivers remarkable performance gains and outperforms all competitive methods significantly across diverse colorization tasks. A direct combination of 3DGS and 2D colorization models leads to significantly higher Matching Error (ME), indicating severe multi-view inconsistency. While ColorNeRF Cheng et al. (2024b) improves 3D color consistency by averaging inconsistent 2D colorizations, this strategy inevitably sacrifices color richness and controllability, resulting in noticeable degradation in FID scores and other task-specific evaluation metrics. We also demonstrate visual comparisons in Fig. 4. As suggested, our method effectively accommodates various forms of user control and accurately generates the desired colorized scenes while preserving strong multi-view consistency. In contrast, direct integration of 2D colorizers introduces obvious view inconsistency and visual artifacts, while ColorNeRF produces overly uniform color tones with limited color diversity and controllability.
@@ 4.3 Results on Dynamic 3D Scene Colorization
The quantitative and qualitative experimental results for the dynamic 3D scenarios are exhibited in the bottom portion of Tab. 1 and Fig. 5. It can be observed that our method significantly outperforms the combination of 4DGS and 2D colorizers, particularly under language-guided and reference-based settings, achieving better FID scores along with substantial reductions in Matching Error (ME) by 0.83 and 0.70, respectively. Moreover, the visual results further demonstrate that our method offers more precise controllability and superior consistency both spatially and temporally.
@@ 4.5 Ablation Studies
    Baseline 115.66 0.6145 0.078
++ Key View Selection 100.28 (+15.38) 0.6175 (+0.0030) 0.074 (+0.004)
++ Single View Augmentation 85.25   (+15.03) 0.6196 (+0.0021) 0.069 (+0.005)
++ Fine-Tuning-Based Colorizer 75.48   (+9.77) 0.6225 (+0.0029) 0.064 (+0.005)
++ Lab Gaussian 68.23   (+7.25) 0.6246 (+0.0021) 0.058 (+0.006)
@@ A.2 More Method Details and Analysis
A.3 More Experiments

### v2

@@ Abstract
In this work, we present Color3D, a highly adaptable framework for colorizing both static and dynamic 3D scenes from monochromatic inputs, delivering visually diverse and chromatically vibrant reconstructions with flexible user-guided control. In contrast to existing methods that focus solely on static scenarios and enforce multi-view consistency by averaging color variations which inevitably sacrifice both chromatic richness and controllability, our approach is able to preserve color diversity and steerability while ensuring cross-view and cross-time consistency. In particular, the core insight of our method is to colorize only a single key view and then fine-tune a personalized colorizer to propagate its color to novel views and time steps. Through personalization, the colorizer learns a scene-specific deterministic color mapping underlying the reference view, enabling it to consistently project corresponding colors to the content in novel views and video frames via its inherent inductive bias. Once trained, the personalized colorizer can be applied to infer consistent chrominance for all other images, enabling direct reconstruction of colorful 3D scenes with a dedicated Lab color space Gaussian splatting representation. The proposed framework ingeniously recasts complicated 3D colorization as a more tractable single image paradigm, allowing seamless integration of arbitrary image colorization models with enhanced flexibility and controllability. Extensive experiments across diverse static and dynamic 3D colorization benchmarks substantiate that our method can deliver more consistent and chromatically rich renderings with precise user control. Project Page: https://yecongwan.github.io/Color3D/.
@@ 1 Introduction
To mitigate the color inconsistency, a few attempts Dhiman et al. (2023); Cheng et al. (2024b) aim to average multi-view color variations via distillation Dhiman et al. (2023) or dynamic color injection Cheng et al. (2024b). While partially effective, they inevitably dilute the palette richness, producing desaturated and tone-flattened results. Furthermore, smoothing color variations makes the colorized results unpredictable, sacrificing user-controlled colorization ability. Moreover, existing studies only focus on static scenes, while controllable colorization of dynamic ones remains an open and unexplored problem, where color consistency in spatial and temporal dimensions should be maintained.
@@ 3.2 Stage 1: Personalized Colorizer Training
Single View Augmentation. In order to avoid overfitting and more robustly generalize to novel views and video frames, we propose a single view augmentation scheme that combines generative augmentations and traditional augmentations to generate more samples in which the same content consistently retains its color across all instances. As illustrated in Fig. 3(a), the proposed scheme first leverages generative models to produce plausible scene content beyond the provided single view, while ensuring chromatic coherence with the input to enrich the diversity of colorization supervision. Specifically, given a single colored view, we incorporate three complementary generative strategies: (1) Outpainting: we divide the image into a $2\times 2$ grid and apply Stable Diffusion Rombach et al. (2022) to each of the four regions individually to imagine plausible scene content beyond their outer boundaries. Meanwhile, we employ LLaVA Liu et al. (2023a) to generate a descriptive caption based on the image, which serves as the text prompt to guide the diffusion process. (2) Image-to-Video: to simulate newly appearing objects and motions in dynamic scenes, we generate continuous video frames conditioned on a single image using Stable Video Diffusion Blattmann et al. (2023). This approach provides coherent object motions and gradual changes in viewpoint, enabling the personalized colorizer to learn more robust color assignments for content that emerges over time. (3) Novel View: to simulate consistent novel viewpoints, we employ Stable Virtual Camera Zhou et al. (2025) to generate views along a predefined orbital trajectory with consistent content across perspectives. Notably, we do not require the generated content to be completely the same as the scene to be colored; instead, we emphasize preserving consistent chromatic styles with the input view, which is sufficient to improve coloring robustness and enhance coloring richness for scene content not included in the single view.
@@ 3.3 Stage 2: 3D Scene Colorization
Table 1: Quantitative comparisons on static (DL3DV-140, LLFF, and Mip-NeRF 360) 3D scene colorization benchmarks. The top-performing results are highlighted with color.
Method Language Automatic Reference
@@
DL3DV-140 (140 Static 3D Scenes)
3DGS+ImageColorizer - - - - 63.56 28.15 0.146 0.038 - - - -
3DGS+VideoColorizer - - - - 77.89 22.38 0.128 0.031 - - - -
Color3D (Ours) - - - - 37.48 32.65 0.084 0.017 - - - -
@@
GaussianEditor 136.49 0.6346 0.093 0.017 - - - - - - - -
GenN2N - - - - 52.63 25.24 0.078 0.015 - - - -
Ref-NPR - - - - - - - - 111.93 0.7958 0.112 0.018
ColorNeRF 138.62 0.6388 0.089 0.018 58.26 21.78 0.062 0.012 103.68 0.7891 0.086 0.013
3DGS+ImageColorizer 87.45 0.6415 0.178 0.035 40.25 34.07 0.084 0.017 67.48 0.7366 0.131 0.023
3DGS+VideoColorizer 118.25 0.6397 0.092 0.017 51.37 26.68 0.069 0.015 88.19 0.7493 0.095 0.015
Color3D (Ours) 62.84 0.6544 0.051 0.009 35.10 33.99 0.056 0.007 41.36 0.6823 0.055 0.008
@@
GaussianEditor 123.71 0.6045 0.088 0.020 - - - - - - - -
GenN2N - - - - 83.69 23.44 0.137 0.032 - - - -
Ref-NPR - - - - - - - - 146.32 0.7928 0.143 0.026
ColorNeRF 163.12 0.6075 0.093 0.020 87.42 18.90 0.093 0.021 121.57 0.7852 0.106 0.021
3DGS+ImageColorizer 112.33 0.6148 0.202 0.042 62.32 29.32 0.135 0.027 86.25 0.7521 0.178 0.032
3DGS+VideoColorizer 119.24 0.6122 0.104 0.023 82.17 23.69 0.099 0.022 92.47 0.7582 0.122 0.024
Color3D (Ours) 68.23 0.6246 0.058 0.012 39.03 33.36 0.082 0.016 48.62 0.7028 0.079 0.015
Table 2: Quantitative comparisons on dynamic (DyNeRF and HyperNeRF) 3D scene colorization benchmarks. The top-performing results are highlighted with color.
Method Language Automatic Reference
[…]
@@
Instruct 4D-to-4D 112.35 0.6082 0.062 0.011 - - - - - - - -
4DGS+ImageColorizer 89.39 0.6159 0.124 0.025 39.17 29.45 0.080 0.018 84.65 0.7446 0.128 0.024
4DGS+VideoColorizer 88.55 0.6199 0.071 0.014 42.36 23.42 0.073 0.014 91.22 0.7598 0.098 0.018
Color3D (Ours) 58.62 0.6271 0.041 0.007 37.28 32.75 0.062 0.009 52.77 0.6717 0.052 0.008
HyperNeRF (Dynamic 3D Scenes)
Instruct 4D-to-4D 118.62 0.6053 0.065 0.012 - - - - - - - -
4DGS+ImageColorizer 105.23 0.6175 0.132 0.029 40.32 29.66 0.090 0.017 76.28 0.7422 0.105 0.018
4DGS+VideoColorizer 110.46 0.6128 0.095 0.017 41.85 24.68 0.084 0.015 89.33 0.7482 0.085 0.014
Color3D (Ours) 63.28 0.6257 0.045 0.008 37.99 33.68 0.067 0.010 58.63 0.6610 0.055 0.008
@@ 4.1 Experimental Settings
Implementation details. All experiments are conducted using the PyTorch framework on NVIDIA RTX A6000 GPUs. We evaluate our method on three static datasets (DL3DV-140 Ling et al. (2024), LLFF Mildenhall et al. (2019) and Mip-NeRF 360 Barron et al. (2022)) and two dynamic datasets (DyNeRF Li et al. (2022) and HyerNeRF Park et al. (2021b)), all of which are converted to monochrome images for colorization. Structure-from-Motion (SfM) is employed to initialize Gaussian points from monochrome inputs in all experiments. Without loss of generality, both our method and competing approaches deploy the same set of image colorization models for fair comparison: ControlColor Liang et al. (2024) for language-guided colorization, DDColor Kang et al. (2023) for automatic colorization, and UniColor Huang et al. (2022) for reference-based colorization. Apart from our proposed Lab Gaussian representation and optimization objectives, all other settings are kept identical to the original 3DGS and 4DGS frameworks. Notably, our entire pipeline incurs only about eight additional minutes for personalized colorizer tuning, which is considered acceptable for chromatic 3D scene reconstruction.
Metrics. We adopt CLIP score Radford et al. (2021) for language-guided colorization to measure text-image alignment, Colorful Score Hasler and Suesstrunk (2003) for automatic colorization to assess color vividness, and Ref-LPIPS Zhang et al. (2023) for reference-based colorization to evaluate perceptual similarity. Fréchet Inception Distance (FID) Heusel et al. (2017) is employed as a standard metric for assessing the rendering quality of all colorization tasks. Additionally, to evaluate the consistency of colorization across views and time, we propose a metric termed Matching Error (ME). It utilizes a dense image matching model Shen et al. (2024) to identify pixel correspondences between ground-truth color images from different views or time steps, and computes the average color difference at these matched locations in the rendered colorized results. Additionally, we compute consistency metrics Dhiman et al. (2023) based on optical-flow correspondences across views and along temporal sequences to more comprehensively assess color consistency across different methods. We report the average of the short-term and long-term consistency scores, which we denote as TC.
@@ 4.2 Results on Static 3D Scene Colorization
The experimental results for quantitative comparisons in static 3D scenes are shown in the middle and upper portions of Tab. 1. Apart from image colorization baselines, we also include a state-of-the-art video colorization model, ColorMNet Yang et al. (2024a), as a reference for more comprehensive evaluations. It is observed that our Color3D delivers remarkable performance gains and outperforms all competitive methods significantly across diverse colorization tasks. A direct combination of 3DGS and image colorization models leads to significantly higher Matching Error (ME), indicating severe multi-view inconsistency. While ColorNeRF Cheng et al. (2024b) improves 3D color consistency by averaging inconsistent 2D colorizations, this strategy inevitably sacrifices color richness and controllability, resulting in noticeable degradation in FID scores and other task-specific evaluation metrics. We also demonstrate visual comparisons in Fig. 4. As suggested, our method effectively accommodates various forms of user control and accurately generates the desired colorized scenes while preserving strong multi-view consistency. In contrast, direct integration of image colorizers introduces obvious view inconsistency and visual artifacts, while ColorNeRF produces overly uniform color tones with limited color diversity and controllability.
@@ 4.3 Results on Dynamic 3D Scene Colorization
The quantitative and qualitative experimental results for the dynamic 3D scenarios are exhibited in the bottom portion of Tab. 1 and Fig. 5. It can be observed that our method significantly outperforms the combination of 4DGS and image colorizers, particularly under language-guided and reference-based settings, achieving better FID scores along with substantial reductions in Matching Error (ME) by 0.83 and 0.70, respectively. Moreover, the visual results further demonstrate that our method offers more precise controllability and superior consistency both spatially and temporally.
@@ 4.4 Results on Real-World Applications
4.5 Diagnostic Analysis of Color Consistency
To better understand the internal behavior of our personalized colorizer and further validate the mechanism behind its strong cross-view consistency, we conducted a diagnostic experiment focusing on the feature representations of the trained model.
Figure 7: Visualization of cross-view feature consistency and color consistency.
Specifically, we analyze whether the colorizer produces consistent and stable features across different viewpoints of the same 3D scene content — a crucial prerequisite for achieving view-consistent color predictions. To this end, we extract features from two distinct camera viewpoints and project them into a low-dimensional space using PCA for visualization. The reduced features are then clustered and color-coded such that similar features share the same color, enabling intuitive comparison across views. As illustrated in Fig. 7, for the same content in 3D space — such as flower clusters, pillars, or stone paths — the extracted features remain consistent and stable across different viewpoints. This indicates that the trained colorizer maintains highly stable feature embeddings for the same 3D content despite changes in camera perspective. Such feature-level consistency directly aligns with the model’s final color outputs and provides strong empirical evidence supporting the design rationale of our personalized colorizer. Please refer to Appendix A.3 for more rigorously theoretical analysis.
@@ 4.6 Ablation Studies
    Baseline 115.66 0.6145 0.078 0.018
++ Key View Selection 100.28 (+15.38) 0.6175 (+0.0030) 0.074 (+0.004) 0.017 (+0.001)
++ Single View Augmentation 85.25   (+15.03) 0.6196 (+0.0021) 0.069 (+0.005) 0.016 (+0.001)
++ Fine-Tuning-Based Colorizer 75.48   (+9.77) 0.6225 (+0.0029) 0.064 (+0.005) 0.013 (+0.003)
++ Lab Gaussian 68.23   (+7.25) 0.6246 (+0.0021) 0.058 (+0.006) 0.012 (+0.001)
@@ A.2 More Method Details and Analysis
A.3 Theoretical analysis of why the learned mapping remains consistent under large view changes.
To help readers gain a deeper understanding of how the proposed personalized colorizer achieves multi-view consistent color propagation, this section provides rigorous convergence guarantees that theoretically establish its effectiveness.
Notation.
Let $\mathcal{S}$ denote the scene surface (luminance). For a 3D surface point $X\in\mathcal{S}$ and a camera/view $v\in\mathcal{V}$, let $I_{v}(X)$ denote the image patch (or local crop) containing the projection of $X$. Let $\phi:\mathcal{I}\to\mathbb{R}^{d}$ be the pretrained encoder (frozen in our pipeline), and $\psi:\mathbb{R}^{d}\to\mathbb{R}^{c}$ be the lightweight adapter / color decoder (which we optimize per scene). Define the composite $F:=\psi\circ\phi$. For two views $v,v^{\prime}$ of the same 3D point $X$ we define the color discrepancy
$\Delta_{c}(X;v,v^{\prime}):=\|F(I_{v}(X))-F(I_{v^{\prime}}(X))\|.$
High-level goal.
Show (under explicit assumptions) (i) a decomposed upper bound on $\Delta_{c}$ that isolates encoder and adapter/decoder contributions, and (ii) how single-view augmentation + consistency training reduces the bound’s dominant terms. Provide (iii) convergence statements for adapter/decoder optimization under clearly stated assumptions.
A.3.1 Assumptions
We explicitly state the assumptions used below. These are empirical/practical but necessary for rigorous statements.
  1. A1
    (Encoder local Lipschitzness) There exists $L_{\phi}>0$ and $\epsilon_{\phi}\geq 0$ such that for any patches $I,I^{\prime}$,
    $\|\phi(I)-\phi(I^{\prime})\|\leq L_{\phi}\|I-I^{\prime}\|+\epsilon_{\phi}.$
    The term $\epsilon_{\phi}$ models finite-capacity / non-smooth residual errors and is assumed small when $I,I^{\prime}$ are semantically similar.
  2. A2
    (Approximate viewpoint equivariance) For typical geometric viewpoint transform $T$ relating views of the same 3D point,
    $\|\phi(I)-\phi(T[I])\|\leq\delta_{\mathrm{eq}},$
    where $\delta_{\mathrm{eq}}$ is small under moderate viewpoint changes (this expresses encoder’s empirical multi-view invariance).
  3. A3
    (Adapter local Lipschitzness) The adapter $\psi$ is locally Lipschitz on the relevant feature domain: there exists $L_{\psi}>0$ such that for features $z,z^{\prime}$,
    $\|\psi(z)-\psi(z^{\prime})\|\leq L_{\psi}\|z-z^{\prime}\|.$
    In practice, $L_{\psi}$ can be controlled via architecture size and weight regularization.
  4. A4
    (Augmentation consistency objective) Training uses an empirical loss combining supervised reference fit (the augmented images) and a consistency term:
    $\mathcal{L}(\theta)\;=\;\mathbb{E}_{X}\Big[\sum_{k}\ell\big(\psi_{\theta}(\phi(T_{k}[I(X)])),c^{\star}(X)\big)\Big]\;+\;\lambda\mathbb{E}_{X}\Big[\sum_{k\neq k^{\prime}}\|\psi_{\theta}(\phi(T_{k}[I]))-\psi_{\theta}(\phi(T_{k^{\prime}}[I]))\|^{2}\Big],$
    where $\theta$ are adapter parameters, $T_{k}$ are single-view augmentations, and $\lambda\geq 0$.
A.3.2 Lemma 1 (Feature difference upper bound).
Under A1–A2, for any two views $v,v^{\prime}$ of the same 3D point $X$,
$\|\phi(I_{v}(X))-\phi(I_{v^{\prime}}(X))\|\leq L_{\phi}\|I_{v}(X)-I_{v^{\prime}}(X)\|+\epsilon_{\phi}+\delta_{\mathrm{eq}}.$
Proof.
By triangle inequality, write $I_{v^{\prime}}\approx T[I_{v}]+\eta$ where $\eta$ models sampling/noise/occlusion residual. Then
$\|\phi(I_{v})-\phi(I_{v^{\prime}})\|\leq\|\phi(I_{v})-\phi(T[I_{v}])\|+\|\phi(T[I_{v}])-\phi(I_{v^{\prime}})\|.$
By A2 the first term is $\leq\delta_{\mathrm{eq}}$. For the second term apply A1 with $I^{\prime}=I_{v^{\prime}}$ and $I=T[I_{v}]$, noting $\|T[I_{v}]-I_{v^{\prime}}\|\leq\|I_{v}-I_{v^{\prime}}\|+\|\eta\|$; absorbing $\|\eta\|$ into $\epsilon_{\phi}$ yields the stated bound. ∎
A.3.3 Lemma 2 (Color difference via feature difference).
Under A3, for any two views $v,v^{\prime}$,
$\Delta_{c}(X;v,v^{\prime})\leq L_{\psi}\cdot\|\phi(I_{v}(X))-\phi(I_{v^{\prime}}(X))\|.$
Proof.
Immediate from the Lipschitz property of $\psi$. ∎
A.3.4 Proposition 1 (Decomposed color discrepancy bound).
Combining Lemmas A.3.2 and A.3.3, for any 3D point $X$ and views $v,v^{\prime}$,
$\boxed{\Delta_{c}(X;v,v^{\prime})\leq L_{\psi}\Big(L_{\phi}\|I_{v}-I_{v^{\prime}}\|+\epsilon_{\phi}+\delta_{\mathrm{eq}}\Big).}$
Proof.
Direct composition of the two lemmas. ∎
Interpretation.
The bound decomposes color difference into: (i) image-level discrepancy $\|I_{v}-I_{v^{\prime}}\|$ (sampling, occlusion), (ii) encoder approximation residual $\epsilon_{\phi}$, and (iii) encoder equivariance error $\delta_{\mathrm{eq}}$. Reducing $L_{\psi}$ (smaller adapter sensitivity) or reducing the effective $\|I_{v}-I_{v^{\prime}}\|$-sensitivity (via consistency training that shrinks feature/ color variance under $T_{k}$) will reduce $\Delta_{c}$.
A.3.5 How consistency training compresses the bound (informal but quantitative sketch).
We formalize the effect of the consistency term by considering the second moment (variance) of color outputs across augmentations for a fixed 3D point:
$V_{\psi}(X):=\mathbb{E}_{T}\big[\|\psi(\phi(T[I(X)]))-m_{\psi}(X)\|^{2}\big],\qquad m_{\psi}(X):=\mathbb{E}_{T}[\psi(\phi(T[I(X)]))].$
The consistency penalty in $\mathcal{L}$ directly bounds $V_{\psi}(X)$. Under standard first-order optimality conditions (or gradient flow dynamics), increasing $\lambda$ reduces the equilibrium value of $V_{\psi}(X)$. Using the Lipschitz relation in Proposition 1 and Jensen, an upper bound on expected squared color discrepancy across sampled view pairs $(v,v^{\prime})$ can be written as
$\mathbb{E}_{X,v,v^{\prime}}[\Delta_{c}^{2}]\leq(L_{\psi}L_{\phi})^{2}\mathbb{E}[\|I_{v}-I_{v^{\prime}}\|^{2}]+(L_{\psi})^{2}(\epsilon_{\phi}+\delta_{\mathrm{eq}})^{2}.$
The consistency term reduces the first RHS term by effectively shrinking the empirical $\mathbb{E}[\|I_{v}-I_{v^{\prime}}\|^{2}]$-sensitivity through learned invariance in $\psi$. We denote this shrinkage by a factor $\eta(\lambda)\in(0,1]$ which empirically decreases as $\lambda$ increases. Thus
$\mathbb{E}[\Delta_{c}^{2}]\leq\eta(\lambda)\,(L_{\psi}L_{\phi})^{2}\mathbb{E}[\|I_{v}-I_{v^{\prime}}\|^{2}]+(L_{\psi})^{2}(\epsilon_{\phi}+\delta_{\mathrm{eq}})^{2}.$
A precise closed-form for $\eta(\lambda)$ depends on dynamics/architecture; we treat $\eta(\lambda)$ as observable and empirically verifiable (see diagnostics below).
A.3.6 Jacobian / sensitivity perspective (formalization).
Let $J_{F}(I):=\nabla_{I}F(I)$ denote the Jacobian of $F=\psi\circ\phi$ wrt image pixels (vectorized). By first-order Taylor expansion, for small image perturbations $\delta I$,
$\|F(I+\delta I)-F(I)\|\leq\|J_{F}(I)\|\cdot\|\delta I\|.$
By chain rule $J_{F}(I)=J_{\psi}(\phi(I))J_{\phi}(I)$ and hence $\|J_{F}\|\leq\|J_{\psi}\|\cdot\|J_{\phi}\|$. Consistency training implicitly penalizes $\|J_{\psi}\|$ (or its expected norm) because reducing output variance across known transforms forces smaller directional derivatives along transform-induced directions; therefore the product $\|J_{F}\|$ is reduced, decreasing sensitivity to viewpoint-induced perturbations.
A.3.7 Convergence results for adapter/decoder optimization
We give convergence statements: Assume $\psi_{\theta}$ is $L$-smooth in parameters (i.e., $\nabla_{\theta}\mathcal{L}$ is Lipschitz with constant $L$) and $\mathcal{L}$ is lower bounded. Then gradient descent (or stochastic gradient descent under standard assumptions) converges to a first-order stationary point: $\lim_{t\to\infty}\|\nabla_{\theta}\mathcal{L}(\theta_{t})\|=0$. If we further assume the Polyak–Łojasiewicz (PL) condition holds locally (i.e., there exists $\mu>0$ with $\tfrac{1}{2}\|\nabla\mathcal{L}(\theta)\|^{2}\geq\mu(\mathcal{L}(\theta)-\mathcal{L}^{\star})$), then gradient descent converges linearly to a global minimizer with rate $1-\alpha\mu$.
Concluding statement.
The preceding derivation has provided (i) a decomposed upper bound on cross-view color discrepancy, (ii) a Jacobian-based mechanism describing how consistency training reduces sensitivity to viewpoint-induced perturbations, and (iii) standard stationary-point results for smooth nonconvex adapter/decoder. The above analysis theoretically substantiates that the proposed personalized colorizer, when optimized with consistency constraints, is guaranteed to produce view-invariant color predictions across varying viewpoints.
A.4 Method Extension
All-in-one training scheme for large-scale datasets. Essentially, our personalized colorizer only includes a scene-specific adapter and a lightweight decoder, while the large encoder is shared across all scenes. This design enables us to jointly train a large number of scenes using a routing-based strategy. The semantic illustration of this mechanism is provided in Fig. 10. Concretely, images from different scenes are stacked along the batch dimension, and each sample is routed through its corresponding adapter and decoder based on the scene ID, which enables efficient all-in-one training. By applying this all-in-one personalized colorizer training scheme, we are able to train colorizers for all 140 scenes in the DL3DV-140 dataset within only 55 minutes. This experiment further demonstrates the practicality and efficiency of our approach for large-scale real-world deployment.
Figure 10: Illustration of the proposed all-in-one training scheme for large-scale datasets.
Generalized to feed-forward style paradigm. To further explore the applicability of the proposed framework beyond per-scene optimization, we investigate its extension to a feed-forward 3D generation paradigm. As illustrated in Fig. 11, we first apply our proposed single-view generative augmentation strategy on a large-scale image dataset to synthesize substantial spatial and viewpoint variations. Using only one colored image as the key view, we train a generalizable personalized colorizer based on a transformer network that learns to colorize all augmented views in a feed-forward manner. At inference time, we follow the exact pipeline described in the paper: we colorize a single input view and then propagate its chroma to other viewpoints using the generalizable personalized colorizer, without any per-scene optimization. The resulting set of multi-view colorized images can then be fed into a feed-forward 3D reconstruction model—such as AnySplat Jiang et al. (2025)—to directly obtain a fully colored Gaussian representation. We trained the generalizable personalized colorizer on the COCO dataset and validated the feed-forward 3D colorization pipeline on the DL3DV-10-Benchmark, demonstrating the feasibility of this extension.
As shown in Tab. 4 and Fig. 12, our experiments indicate that the proposed framework already exhibits promising feed-forward 3D colorization potential. While a fully feed-forward solution is indeed appealing, our observations show that a globally generalized colorizer still falls short of our per-scene adapter in both chromatic richness and multi-view consistency, and it suffers from significant resolution degradation. Given the large-baseline viewpoints and high-resolution reconstruction settings considered in our work, the per-scene optimization route remains the most practical and reliable choice at present. Nonetheless, this preliminary investigation highlights a viable direction for future research.
Figure 11: Illustration of the proposed feed-forward style 3D colorization framework.
Table 4: Quantitative comparisons of automatic colorization on the DL3DV-140 dataset.
[…]
Generalizable Colorizer + AnySplat 82.36 24.37 0.142 0.030
Per-Scene Colorizer + AnySplat 66.78 29.39 0.109 0.025
Figure 12: Visual comparison of generalizable personalized colorizer and per-scene personalized colorizer for feed-forward 3D colorization.
A.5 More Experiments